Sir/ma'am,
Please explain the sequence to me
2,7,5,9,9,1,14,13,---,---
answer is 20,15
thanks

Answer by ramkikk66(644)


Your series is: 2,7,5,9,9,1,14,13,---,--- 

Pl check - is the 6th term 1 or 11?

I think there could be a typo and it is actually 11.

If so, the series is 2,7,5,9,9,11,14,13,...,... and it can be solved.

Solution:

Look at alternate terms - they form 2 separate series. 

Series 1:

2,5,9,14... where the difference between successive terms keeps increasing by 1. 
2 + 3 = 5
5 + 4 = 9
9 + 5 = 14
So the next term in this series is 14 + 6 = 20

Series 2:
And the other series is
7,9,13,... which is a simple arithmetic progression with difference as 2. So the next term in this series is 13 + 2 = 15.

So the next two terms are 20 and  15.

Hope this helps.

:)